what is 3x-4(5+x)=x+8
Answer by scoop(27)

use the distribution method
3x-20-4x=x+8
combine like terms
3x-4x-20=x+8
-x-20=x+8
get variables on one side
-20=2x+8
-20-8=2x
-28=2x
divide both sides by 2
x=-14
if you plug -14 in for all of your x variables you will see that it checks